### v2

@@ 1 Introduction
However, two major difficulties currently impede operator inference in accurately reconstructing the intrusive ROM operators. First, the FOM snapshot data contains non-Markovian information that is inconsistent with the Markovian intrusive ROM. This non-Markovian information perturbes the operator inference minimization problem such that the intrusive ROM generally does not minimize this perturbed problem. To eliminate this perturbation, Peherstorfer has proposed to generate additional FOM snapshot data whereby the FOM state is projected onto the ROM basis after each time step [18], known as the re-projection approach. The resulting FOM snapshot data does not contain non-Markovian information and is equivalent to snapshot data generated by the intrusive ROM. Consequently, the minimization problem is unperturbed and the intrusive ROM does minimize it. However, operator inference is only guaranteed to yield this solution if this minimizer is unique. Concretely, the operator inference minimization problem is a least squares problem and has a unique solution only if the least squares matrix has full rank.
@@
In this work, we propose a novel snapshot data generation method that guarantees the full rank of the operator inference least squares problem. The number of generated snapshots is equal to this rank and hence minimal. Like Peherstorfer’s approach [18], our method avoids non-Markovian information and hence reconstructs the intrusive ROM operators exactly, i.e., up to machine precision. Our method has two key ingredients. First, we leverage the knowledge about the specific polynomial structure of the intrusive ROM to identify a set of ROM states and input signals that define snapshots that induce a full rank least squares problem. Second, we generate snapshots corresponding to these identified ROM states and input signals by simulating multiple trajectories, initialized with the identified ROM states as initial conditions and the identified input signals, for only a single time step. In contrast to the usual approach that generates snapshot data along long trajectories, this single-step approach gives us control over which snapshot data we obtain. Additionally, this single-step approach avoids any non-Markovian terms, so we do not need re-projection[18]. Since the number of snapshots is minimal, the least squares problem simplifies to a linear system that is numerically more stable. Since the method exactly reconstructs intrusive ROM operators, it preserves structures such as symmetry and skew-symmetry of these intrusive operators [23] without the need to enforce this preservation explicitly [12, 8, 7]. Compared with the re-projection approach [18], our proposed approach is less intrusive because it does not require a re-projection step after each time step.
@@ 2.1 Dynamical systems with polynomial terms
with state ${\bm{x}}(t;{\bm{\mu}})\in\mathbb{R}^{N}$, input ${\bm{u}}(t;{\bm{\mu}})\in\mathbb{R}^{N_{u}}$, time $t\in[0,T]$, and parameter ${\bm{\mu}}\in\mathcal{D}\subset\mathbb{R}^{d}$.
We assume the nonlinear function ${\bm{f}}:\mathbb{R}^{N}\times\mathbb{R}^{N_{u}}\times\mathcal{D}\rightarrow\mathbb{R}^{N}$ to be polynomial of order $l\in\mathbb{N}_{0}$ in the state ${\bm{x}}$ and linear in the input signal ${\bm{u}}$, so there exist ${\mathbf{A}}_{i}({\bm{\mu}})\in\mathbb{R}^{N\times N_{i}}$ with $N_{i}=\begin{pmatrix}N+i-1\\ i\end{pmatrix}$ for $i\in\mathcal{I}$, and ${\mathbf{B}}({\bm{\mu}})\in\mathbb{R}^{N\times{N_{u}}}$ for ${\bm{\mu}}\in\mathcal{D}$ such that
@@
with the degree set $\mathcal{I}\subset\{0,1,\dots,l\}$.
The vector ${\bm{x}}^{i}\in\mathbb{R}^{N_{i}}$ comprises all unique products of $i$ entries of ${\bm{x}}$ (${\bm{x}}^{0}$ is scalar 1). We can construct this vector from the $i$-fold Kronecker product ${\bm{x}}\otimes\dots\otimes{\bm{x}}=:{\bm{x}}^{i_{\otimes}}$ by removing all duplicate entries due to commutativity of the multiplication [19]. Formally, we can write ${\bm{x}}^{i}={\mathbf{I}}_{N}^{(i)}{\bm{x}}^{i_{\otimes}}$ with ${\mathbf{I}}_{N}^{(i)}\in\{0,1\}^{N_{i}\times N^{i}}$ an identity matrix with certain rows removed, and conversely ${\bm{x}}^{i_{\otimes}}={\mathbf{J}}_{N}^{(i)}{\bm{x}}^{i}$ with ${\mathbf{J}}_{N}^{(i)}\in\{0,1\}^{N^{i}\times N_{i}}$ an identity matrix with certain rows copied and permuted. Here $N^{i}$ denotes as usual the $i$-th power of $N$.
@@ 2.2 Intrusive projection-based model reduction
Here, ${\mathbf{V}}_{n}=[{\bm{v}}_{1},\dots,{\bm{v}}_{n}]\in\mathbb{R}^{N\times n}$ is an orthonormal basis of an $n$-dimensional subspace of $\mathbb{R}^{N}$ with $n\ll N$, and $\tilde{\bm{x}}(t;{\bm{\mu}})\in\mathbb{R}^{n}$ the ROM state.
The orthonormal basis ${\mathbf{V}}_{n}$ can, for example, be obtained from proper orthogonal decomposition (POD) [11] of snapshots ${\bm{x}}_{k}({\bm{\mu}})\in\mathbb{R}^{N},\,k=0,\dots,K_{\text{POD}}-1$ of a time discretization of the FOM (1). Other methods to construct reduced bases and details on how to sample parameters and the ROM dimension $n$ are discussed in [4].
@@
whose dimensions only depend on the ROM dimension $n$ (and the unreduced input signal dimension ${N_{u}}$) [18]. Note that these precomputations must in general be performed for each choice of the parameter ${\bm{\mu}}\in\mathcal{D}$ separately.
The resulting efficient ROM for fixed ${\bm{\mu}}$ is given by
@@ 2.3 Operator inference
As described in the previous subsection, intrusive model reduction requires access to the FOM operators ${\mathbf{A}}_{i}({\bm{\mu}})$, ${\mathbf{B}}({\bm{\mu}})$ to precompute the ROM operators $\tilde{\mathbf{A}}_{i}({\bm{\mu}})$, $\tilde{\mathbf{B}}({\bm{\mu}})$ via (4) and (5). In practice, this access is not possible in many FOM solvers. Instead, operator inference aims to learn approximations $\hat{\mathbf{A}}_{i}({\bm{\mu}})$, $\hat{\mathbf{B}}({\bm{\mu}})$ of the intrusive ROM operators from data [19].
This data consists of triples of a ROM state snapshot $\breve{\bm{x}}_{k}({\bm{\mu}})\in\mathbb{R}^{n}$, a ROM state time derivative snapshot $\dot{\breve{\bm{x}}}_{k}({\bm{\mu}})\in\mathbb{R}^{n}$ and an input signal snapshot ${\bm{u}}_{k}({\bm{\mu}})$, $k=0,\dots,K-1$. For these triples, operator inference minimizes the residual of the ROM ODE (6) to infer the aggregated ROM operator,
@@
If snapshot triples are obtained from the intrusive ROM ODE (6) such that $\dot{\breve{\bm{x}}}_{k}({\bm{\mu}})=\dot{\tilde{\bm{x}}}(t_{k};{\bm{\mu}})$, $\breve{\bm{x}}_{k}({\bm{\mu}})=\tilde{\bm{x}}(t_{k};{\bm{\mu}})$ and ${\bm{u}}_{k}({\bm{\mu}})={\bm{u}}(t_{k};{\bm{\mu}})$ at time steps $t_{k}$, $k=0,\dots,K-1$, then the intrusive aggregated ROM operator $\tilde{\mathbf{O}}({\bm{\mu}})=\left[[\tilde{\mathbf{A}}_{i}]_{i\in\mathcal{I}}\;\;\tilde{\mathbf{B}}({\bm{\mu}})\right]$ is a solution to this minimization problem.
@@ 3.1 Rank-ensuring ROM states and input signals
If =$\mathcal{I}=\{1\}$ and ${N_{u}}=0$, so ${\bm{p}}(\tilde{\bm{x}})={\bm{x}}$, we can achieve such an isolation of entries by using the unit vectors in $\mathbb{R}^{n}$, here denoted by ${\bm{e}}_{j}$, $j=1,\dots,n$, as inputs $\tilde{\bm{x}}$.
If =$\mathcal{I}=\{2\}$ and ${N_{u}}=0$, so ${\bm{p}}(\tilde{\bm{x}})=\tilde{\bm{x}}^{2}$, perfect isolation of entries is not possible. For example, when choosing $\tilde{\bm{x}}={\bm{e}}_{i}+{\bm{e}}_{j}$ such that ${\bm{p}}$ is 1 in the entry $x_{i}x_{j}$ with $i\neq j$, also the entries $x_{i}^{2}$ and $x_{j}^{2}$ are 1. However, the columns of $\hat{\mathbf{O}}$ corresponding to $x_{i}x_{j}$, $x_{i}^{2}$ and $x_{j}^{2}$ can still be inferred when employing the data $\hat{\mathbf{O}}{\bm{p}}({\bm{e}}_{i}+{\bm{e}}_{j})$, $\hat{\mathbf{O}}{\bm{p}}(2{\bm{e}}_{i})$ and $\hat{\mathbf{O}}{\bm{p}}(2{\bm{e}}_{j})$.
@@
[…]
@@
As an example, for $n=2$, ${N_{u}}=2$ and =$\mathcal{I}=\{1,2\}$ (such that ${n_{p}}=5$, ${n_{f}}=7$), we get the pairs
@@
${\mathbf{P}}=\begin{bmatrix}1&0&2&1&0&0&0\\ 0&1&0&1&2&0&0\\ 1&0&4&1&0&0&0\\ 0&0&0&1&0&0&0\\ 0&1&0&1&4&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\\ \end{bmatrix}\hskip 10.00002pt\text{ where }\hskip 10.00002pt{\bm{p}}(\tilde{\bm{x}},{\bm{u}})=\begin{bmatrix}x_{1}\\ x_{2}\\ x_{1}^{2}\\ x_{1}x_{2}\\ x_{2}^{2}\\ u_{1}\\ u_{2}\end{bmatrix}.$
A proof that the choice of ROM states and input signals (14) results in full rank of the matrix ${\mathbf{P}}$ for any $n\in\mathbb{N}^{+}$, ${N_{u}}\in\mathbb{N}_{0}$ and $\mathcal{I}\subset\mathbb{N}_{0}$ is given in Appendix A.
Note that this choice of pairs results in a matrix ${\mathbf{P}}$ that has as many columns as rows. This means that our choice uses the minimal number of snapshots necessary to achieve full rank. In addition, because ${\mathbf{P}}$ is square, the operator inference minimization problem (11) is equivalent to the matrix-valued linear system
@@
corresponding to the least squares problem (11), this linear system is in general less sensitive to error amplifications.
@@ 3.2 Snapshot data generation via single-step ensemble
and integrated in time for only a single explicit Euler step of size $\Delta t$. This approach results in FOM snapshots
@@
[…]
The resulting triples $(\dot{\bar{\bm{x}}}_{s},\bar{\bm{x}}_{s},{\bm{u}}_{s})_{s=1}^{n_{f}}$ form the input to the operator inference optimization problem (8). Since the matrix ${\mathbf{P}}$ (12) only depends on the snapshots $\bar{\bm{x}}_{s}$ and ${\bm{u}}_{s}$, $s=1,\dots,{n_{f}}$, it does not depend on the unknown FOM operators. Hence, we can guarantee full rank of ${\mathbf{P}}$ by the choice of the initial conditions and input signals.
Note that our novel approach for generating snapshot data for operator inference does not affect the snapshot data used for POD, so the POD basis ${\mathbf{V}}_{n}$ remains unchanged.
The idea to generate snapshot data in multiple trajectories, each containing only a single time step, has already been used in numerical experiments in [19]. However, randomized instead of well-chosen initial conditions are used in that work. Consequently far more than ${n_{f}}$ trajectories are necessary to achieve full rank of ${\mathbf{P}}$.
@@
5:     for $\bar{\bm{x}}\in\cup_{i\in\mathcal{I}}\mathcal{X}^{i}$ do
6:         ${\bm{x}}_{1}\leftarrow$ FOM( initial condition ${\bm{x}}_{0}={\mathbf{V}}_{n}\bar{\bm{x}}$,
@@
7:         ${\mathbf{P}}\leftarrow[{\mathbf{P}},\;\;{\bm{p}}(\tilde{\bm{x}}_{0},{\bm{0}})]$
8:         $\dot{\breve{{\mathbf{X}}}}\leftarrow[\dot{\breve{{\mathbf{X}}}},\;\;\frac{{\mathbf{V}}_{n}^{T}{\bm{x}}_{1}-\bar{\bm{x}}}{\Delta t}]$
9:     end for
10:     for $j=1,\dots,{N_{u}}$ do
11:         ${\bm{u}}\leftarrow$ $j$-th unit vector in $\mathbb{R}^{N_{u}}$
12:         ${\bm{x}}_{1}\leftarrow$ FOM( initial condition ${\bm{x}}_{0}={\bm{0}}$,
@@
13:         ${\mathbf{P}}\leftarrow[{\mathbf{P}},\;\;{\bm{p}}({\bm{0}},{\bm{u}})]$
14:         $\dot{\breve{{\mathbf{X}}}}\leftarrow[\dot{\breve{{\mathbf{X}}}},\;\;\frac{{\mathbf{V}}_{n}^{T}{\bm{x}}_{1}-{\bm{0}}}{\Delta t}]$
15:     end for
16:     Solve linear system (15)
return $\hat{\mathbf{O}}$
17:end procedure

### v3

@@ 1 Introduction
However, two major difficulties currently impede operator inference in accurately reconstructing the intrusive ROM operators. First, the FOM snapshot data contains non-Markovian information that is inconsistent with the Markovian intrusive ROM. This non-Markovian information perturbs the operator inference minimization problem such that the intrusive ROM generally does not minimize this perturbed problem. To eliminate this perturbation, Peherstorfer has proposed to generate additional FOM snapshot data whereby the FOM state is projected onto the ROM basis after each time step [18], known as the re-projection approach. The resulting FOM snapshot data does not contain non-Markovian information and is equivalent to snapshot data generated by the intrusive ROM. Consequently, the minimization problem is unperturbed and the intrusive ROM does minimize it. However, operator inference is only guaranteed to yield this solution if this minimizer is unique. Concretely, the operator inference minimization problem is a least squares problem and has a unique solution only if the least squares matrix has full rank.
@@
In this work, we propose a novel snapshot data generation method that guarantees the full rank of the operator inference least squares problem. The number of generated snapshots is equal to this rank and hence minimal. Like Peherstorfer’s approach [18], our method avoids non-Markovian information and hence reconstructs the intrusive ROM operators exactly, i.e., up to machine precision. Our method has two key ingredients. First, we leverage the knowledge about the specific polynomial structure of the intrusive ROM to identify a set of ROM states and input signals that define snapshots that induce a full rank least squares problem. Second, we generate snapshots corresponding to these identified ROM states and input signals by simulating multiple trajectories, initialized with the identified ROM states as initial conditions and the identified input signals, for only a single explicit Euler time step. In contrast to the usual approach that generates snapshot data along long trajectories, this single-step approach gives us control over which snapshot data we obtain. Additionally, this single-step approach avoids any non-Markovian terms, so we do not need re-projection[18], and the explicit Euler scheme prevents time discretization inconsistencies. Since the number of snapshots is minimal, the least squares problem simplifies to a linear system that is numerically more stable. Since the method exactly reconstructs intrusive ROM operators, it preserves structures such as symmetry and skew-symmetry of these intrusive operators [24] without the need to enforce this preservation explicitly [12, 8, 7].
@@ 2.1 Dynamical systems with polynomial terms
with state ${\bm{x}}(t;{\bm{\mu}})\in\mathbb{R}^{N}$, input ${\bm{u}}(t;{\bm{\mu}})\in\mathbb{R}^{N_{u}}$, time $t\in[0,T]$, and parameter ${\bm{\mu}}\in\mathcal{D}\subset\mathbb{R}^{d}$. We assume the nonlinear function ${\bm{f}}:\mathbb{R}^{N}\times\mathbb{R}^{N_{u}}\times\mathcal{D}\rightarrow\mathbb{R}^{N}$ to be polynomial of order $l\in\mathbb{N}_{0}$ in the state $\bm{x}$} and linear in the input signal $\bm{u}$}, so there exist ${\mathbf{A}}_{i}({\bm{\mu}})\in\mathbb{R}^{N\times N_{i}}$ with $N_{i}=\begin{pmatrix}N+i-1\\ i\end{pmatrix}$ for $i\in\mathcal{I}$, and ${\mathbf{B}}({\bm{\mu}})\in\mathbb{R}^{N\times{N_{u}}}$ for ${\bm{\mu}}\in\mathcal{D}$ such that
@@
with the degree set $\mathcal{I}\subset\{0,1,\dots,l\}$. The vector ${\bm{x}}^{i}\in\mathbb{R}^{N_{i}}$ comprises all unique products of $i$ entries of $\bm{x}$} (${\bm{x}}^{0}$ is scalar 1). We can construct this vector from the $i$-fold Kronecker product ${\bm{x}}\otimes\dots\otimes{\bm{x}}=:{\bm{x}}^{i_{\otimes}}$ by removing all duplicate entries due to commutativity of the multiplication [19]. Formally, we can write ${\bm{x}}^{i}={\mathbf{I}}_{N}^{(i)}{\bm{x}}^{i_{\otimes}}$ with ${\mathbf{I}}_{N}^{(i)}\in\{0,1\}^{N_{i}\times N^{i}}$ an identity matrix with certain rows removed, and conversely ${\bm{x}}^{i_{\otimes}}={\mathbf{J}}_{N}^{(i)}{\bm{x}}^{i}$ with ${\mathbf{J}}_{N}^{(i)}\in\{0,1\}^{N^{i}\times N_{i}}$ an identity matrix with certain rows copied and permuted. Here $N^{i}$ denotes as usual the $i$-th power of $N$.
@@ 2.2 Intrusive projection-based model reduction
Here, ${\mathbf{V}}_{n}=[{\bm{v}}_{1},\dots,{\bm{v}}_{n}]\in\mathbb{R}^{N\times n}$ is an orthonormal basis of an $n$-dimensional subspace of $\mathbb{R}^{N}$ with $n\ll N$, and $\tilde{\bm{x}}(t;{\bm{\mu}})\in\mathbb{R}^{n}$ the ROM state. The orthonormal basis ${\mathbf{V}}_{n}$ can, for example, be obtained from proper orthogonal decomposition (POD) [11] of snapshots ${\bm{x}}_{k}({\bm{\mu}})\in\mathbb{R}^{N},\,k=0,\dots,K_{\text{POD}}-1$ of a time discretization of the FOM (1). Other methods to construct reduced bases and details on how to sample parameters and the ROM dimension $n$ are discussed in [4].
@@
whose dimensions only depend on the ROM dimension $n$ (and the unreduced input signal dimension $N_{u}$}) [18]. Note that these precomputations must in general be performed for each choice of the parameter ${\bm{\mu}}\in\mathcal{D}$ separately. The resulting efficient ROM for fixed $\bm{\mu}$} is given by
@@ 2.3 Operator inference
As described in the previous subsection, intrusive model reduction requires access to the FOM operators ${\mathbf{A}}_{i}({\bm{\mu}})$, ${\mathbf{B}}({\bm{\mu}})$ to precompute the ROM operators $\tilde{\mathbf{A}}_{i}({\bm{\mu}})$, $\tilde{\mathbf{B}}({\bm{\mu}})$ via (4) and (5). In practice, this access is not possible in many FOM solvers. Instead, operator inference aims to learn approximations $\hat{\mathbf{A}}_{i}({\bm{\mu}}){\;\in\mathbb{R}^{n\times n_{i}}}$, $\hat{\mathbf{B}}({\bm{\mu}}){\;\in\mathbb{R}^{n\times{N_{u}}}}$ of the intrusive ROM operators from data [19]. This data consists of triples of a ROM state snapshot $\breve{\bm{x}}_{k}({\bm{\mu}})\in\mathbb{R}^{n}$, a ROM state time derivative snapshot $\dot{\breve{\bm{x}}}_{k}({\bm{\mu}})\in\mathbb{R}^{n}$ and an input signal snapshot ${\bm{u}}_{k}({\bm{\mu}})$, $k=0,\dots,K-1$. For these triples, operator inference minimizes the residual of the ROM ODE (6) to infer the aggregated ROM operator,
@@
If snapshot triples are obtained from the intrusive ROM ODE (6) such that $\breve{\bm{x}}_{k}({\bm{\mu}})=\tilde{\bm{x}}(t_{k};{\bm{\mu}})$, $\dot{\breve{\bm{x}}}_{k}({\bm{\mu}})=\dot{\tilde{\bm{x}}}(t_{k};{\bm{\mu}})$ and ${\bm{u}}_{k}({\bm{\mu}})={\bm{u}}(t_{k};{\bm{\mu}})$ at time steps $t_{k}$, $k=0,\dots,K-1$, then the intrusive aggregated ROM operator $\tilde{\mathbf{O}}({\bm{\mu}})=\left[[\tilde{\mathbf{A}}_{i}]_{i\in\mathcal{I}}\;\;\tilde{\mathbf{B}}({\bm{\mu}})\right]$ is a solution to this minimization problem.
@@ 3.1 Rank-ensuring ROM states and input signals
Example 1.
If $\mathcal{I}=\{1\}$ and ${N_{u}}=0$, so ${\bm{p}}(\tilde{\bm{x}})={\tilde{\bm{x}}}$, we can achieve such an isolation of entries by using the unit vectors in $\mathbb{R}^{n}$, here denoted by ${\bm{e}}_{j}$, $j=1,\dots,n$, as inputs $\tilde{\bm{x}}$.
Example 2.
If $\mathcal{I}=\{2\}$ and ${N_{u}}=0$, so ${\bm{p}}(\tilde{\bm{x}})=\tilde{\bm{x}}^{2}$, perfect isolation of entries is not possible. For example, when choosing $\tilde{\bm{x}}\;{=[\tilde{x}_{1},\dots,\tilde{x}_{n}]^{T}}={\bm{e}}_{i}+{\bm{e}}_{j}$ such that ${\bm{p}}{(\tilde{\bm{x}})}$ is 1 in the entry $\tilde{x}_{i}\tilde{x}_{j}$} with $i\neq j$, also the entries $\tilde{x}_{i}^{2}$} and $\tilde{x}_{j}^{2}$} are 1. However, the columns of $\hat{\mathbf{O}}$ corresponding to $\tilde{x}_{i}\tilde{x}_{j}$}, $\tilde{x}_{i}^{2}$} and $\tilde{x}_{j}^{2}$} can still be inferred when employing the data $\hat{\mathbf{O}}{\bm{p}}({\bm{e}}_{i}+{\bm{e}}_{j})$, $\hat{\mathbf{O}}{\bm{p}}(2{\bm{e}}_{i})$ and $\hat{\mathbf{O}}{\bm{p}}(2{\bm{e}}_{j})$.
@@
(13) $\mathcal{X}^{i}:=\left\{\sum_{j=1}^{i}{\bm{a}}_{j}\Bigg|{\bm{a}}_{1},\dots,{\bm{a}}_{i}\in\{{\bm{e}}_{1},\dots,{\bm{e}}_{n}\}\right\}\hskip 10.00002pt\text{ and consistently }\mathcal{X}^{0}:=\{{\bm{0}}\in\mathbb{R}^{n}\}.$
@@
Example 3.
For $n=2$, ${N_{u}}=2$ and $\mathcal{I}=\{1,2\}$ (such that ${n_{p}}=5$, ${n_{f}}=7$), we get the pairs
@@
${\mathbf{P}}=\begin{bmatrix}1&0&2&1&0&0&0\\ 0&1&0&1&2&0&0\\ 1&0&4&1&0&0&0\\ 0&0&0&1&0&0&0\\ 0&1&0&1&4&0&0\\ 0&0&0&0&0&1&0\\ 0&0&0&0&0&0&1\\ \end{bmatrix}\hskip 10.00002pt\text{ where }\hskip 10.00002pt{\bm{p}}(\tilde{\bm{x}},{\bm{u}})={\begin{bmatrix}\tilde{x}_{1}\\ \tilde{x}_{2}\\ \tilde{x}_{1}^{2}\\ \tilde{x}_{1}\tilde{x}_{2}\\ \tilde{x}_{2}^{2}\\ u_{1}\\ u_{2}\end{bmatrix}}$.
In general, the following theorem holds.
[…]
For any $n\in\mathbb{N}^{+}$, ${N_{u}}\in\mathbb{N}_{0}$ and $\mathcal{I}\subset\mathbb{N}_{0}$, the matrix ${\mathbf{P}}:=\left[{\bm{p}}(\breve{\bm{x}}_{0},{\bm{u}}_{0})\;\;\dots\;\;{\bm{p}}(\breve{\bm{x}}_{K-1},{\bm{u}}_{K-1})\right]\;\in\mathbb{R}^{{n_{f}}\times K}$ defined in (12) with $K={n_{f}}$, ${\bm{p}}(\tilde{\bm{x}},{\bm{u}})=\begin{bmatrix}\left[\tilde{\bm{x}}^{i}\right]_{i\in\mathcal{I}}\\ {\bm{u}}\end{bmatrix}$ as defined in (7) and with the ROM states and input signals $\left\{(\bar{\bm{x}}_{j},\bar{\bm{u}}_{j})\right\}_{j=1}^{n_{f}}=\left\{(\bar{\bm{x}}_{j},{\bm{0}})|\tilde{\bm{x}}_{j}\in\cup_{i\in\mathcal{I}}\mathcal{X}^{i}\right\}\cup\left\{({\bm{0}},{\bm{u}}_{j})\right\}_{j=1}^{N_{u}}$ defined in (14) has full rank.
The proof is given in Appendix A.
Note that this choice of ROM states and input signals results in a matrix $\mathbf{P}$} that has as many columns as rows. This means that our choice uses the minimal number of snapshots necessary to achieve full rank. In addition, because $\mathbf{P}$} is square, the operator inference minimization problem (11) is equivalent to the matrix-valued linear system
@@
corresponding to the least squares problem (11), the linear system (15) is in general less sensitive to error amplifications.
@@ 3.2 Snapshot data generation via single-step ensemble
with $\bar{\bm{x}}_{s}$ and $\bar{\bm{u}}_{s}$ as defined in (14), and integrated in time for only a single explicit Euler step of size $\Delta t$. This approach results in FOM snapshots
@@
(19) $\dot{\bar{\bm{x}}}_{s}={\mathbf{V}}_{n}^{T}\frac{{\bm{x}}_{1}^{s}-{\bm{x}}_{0}^{s}}{\Delta t}\hskip 10.00002pt\hskip 10.00002pt\hskip 10.00002pts=1,\dots,{n_{f}}.$
The resulting triples $({\bar{\bm{x}}_{s},\dot{\bar{\bm{x}}}_{s}},{\bm{u}}_{s})_{s=1}^{n_{f}}$ form the input to the operator inference optimization problem (8). Since the matrix $\mathbf{P}$} (12) only depends on the snapshots $\bar{\bm{x}}_{s}$ and ${\bm{u}}_{s}$, $s=1,\dots,{n_{f}}$, it does not depend on the unknown FOM operators. Hence, we can guarantee full rank of $\mathbf{P}$} by the choice of the initial conditions and input signals.
Note that our novel approach for generating snapshot data for operator inference does not affect the snapshot data used for POD, so the POD basis ${\mathbf{V}}_{n}$ remains unchanged. For example, the implicit Euler method is used to generate POD snapshot data in Section 4.2.
The idea to generate snapshot data in multiple trajectories, each containing only a single time step, has already been used in numerical experiments in [19]. However, randomized instead of well-chosen initial conditions are used in that work. Later, FOM snapshots projected onto the ROM space have been used as initial conditions in context of lifted nonlinearities [21] and noisy data [29]. All three approaches enable to check the full rank a priori, but do not guarantee full rank. Indeed, far more than $n_{f}$} trajectories have been used to achieve full rank of $\mathbf{P}$} in the numerical experiments in these works.
@@
5:  for $i\in\mathcal{I}$ do
6:   # Generate ROM state sets $\mathcal{X}^{i}$ as defined in (13) (see Appendix B)
7:   $\mathcal{X}^{i}\leftarrow$ GenerateROMstateSet(ROM dimension $n$, polynomial degree ii)
8:  end for
9:   for $\bar{\bm{x}}\in\cup_{i\in\mathcal{I}}\mathcal{X}^{i}$ do
10:     ${\bm{x}}_{1}\leftarrow$ FOM( initial condition ${\bm{x}}_{0}={\mathbf{V}}_{n}\bar{\bm{x}}$,
@@
11:     ${\mathbf{P}}\leftarrow[{\mathbf{P}},\;\;{\bm{p}}(\tilde{\bm{x}}_{0},{\bm{0}})]$
12:     $\dot{\breve{{\mathbf{X}}}}\leftarrow[\dot{\breve{{\mathbf{X}}}},\;\;\frac{{\mathbf{V}}_{n}^{T}{\bm{x}}_{1}-\bar{\bm{x}}}{\Delta t}]$
13:   end for
14:   for $j=1,\dots,{N_{u}}$ do
15:     $\bm{u}$$\leftarrow$ $j$-th unit vector in $\mathbb{R}^{N_{u}}$
16:     ${\bm{x}}_{1}\leftarrow$ FOM( initial condition ${\bm{x}}_{0}={\bm{0}}$,
@@
17:     ${\mathbf{P}}\leftarrow[{\mathbf{P}},\;\;{\bm{p}}({\bm{0}},{\bm{u}})]$
18:     $\dot{\breve{{\mathbf{X}}}}\leftarrow[\dot{\breve{{\mathbf{X}}}},\;\;\frac{{\mathbf{V}}_{n}^{T}{\bm{x}}_{1}-{\bm{0}}}{\Delta t}]$
19:   end for
20:   Solve linear system (15)
21:   return $\hat{\mathbf{O}}$
22:  end procedure